log(4+3^5x)=log(8)
Answer by stanbon(48535)

log(4+3^5x)=log(8)
log(4+3^5x)-log(8) = 0

---
log[(4+3^(5x))/8] = 0
---
(4+3^(5x))/8 = 10^0
---
4+3^(5x) = 8
---
3^(5x) = 4
---
5x*log(3) = log(4)
5x = [log(4)]/[log(3)]
---
x = (1/5)[log(4)]/[log(3)]
----
x = 0.2524
=================
Cheers,
Stan H.